# Semantic Regexes: Auto-Interpreting LLM Features with a Structured Language

### v1

@@ Abstract
Automated interpretability aims to translate large language model (LLM) features into human understandable descriptions. However, these natural language feature descriptions are often vague, inconsistent, and require manual relabeling. In response, we introduce semantic regexes, structured language descriptions of LLM features. By combining primitives that capture linguistic and semantic feature patterns with modifiers for contextualization, composition, and quantification, semantic regexes produce precise and expressive feature descriptions. Across quantitative benchmarks and qualitative analyses, we find that semantic regexes match the accuracy of natural language while yielding more concise and consistent feature descriptions. Moreover, their inherent structure affords new types of analyses, including quantifying feature complexity across layers, scaling automated interpretability from insights into individual features to model-wide patterns. Finally, in user studies, we find that semantic regex descriptions help people build accurate mental models of LLM feature activations.
@@ 1 Introduction
AI models represent their learned concepts, like the “Golden Gate Bridge”, as linear directions in latent space, called features (Bricken et al., 2023). Understanding a model’s features helps us anticipate its behavior, assess its human-alignment, and intervene to ensure safe outcomes (Templeton et al., 2024). Towards this goal, automated interpretability assigns human-readable descriptions to each feature by analyzing patterns in its response to input data (Paulo et al., 2024; Bills et al., 2023; Lin, 2023). With feature descriptions, researchers have identified how models encode domain-relevant concepts, like protein structures (Gujral et al., 2025), and reconstructed circuits of features that correspond to complex behaviors, like medical diagnoses (Lindsey et al., 2025).
@@
In contrast, structured languages offer well-defined syntax and semantics (Chomsky, 1956). By combining a constrained set of primitives with compositional rules, structured languages (e.g., regular language and programming languages) can capture precise patterns while maintaining expressivity (Lawson, 2005). Their grammatical structure provides additional affordances, including consistent ways of expressing the same pattern (Knuth, 1965), the ability to concisely capture complex patterns (Lawson, 2005), and mechanisms for representing abstraction (Backus, 1959; Aho & Ullman, 1972). As a result, structured languages can construct complex yet specific expressions, ranging from regular expressions (Lawson, 2005) to software (Python Software Foundation, 2025).
@@
Across automated interpretability evaluations, we find that semantic regexes as as accurate as natural language, showing that constraining feature descriptions to the semantic regex language does not reduce expressivity. Beyond accuracy, their structured form offers additional affordances for interpretability, including producing more concise feature descriptions and enforcing description consistency across functionally similar features. Moreover, since semantic regex primitives and modifiers exist across levels of abstraction, they serve as a proxy for feature complexity, allowing us to perform model-wise analysis of feature behavior. Finally, in a user study, we find that semantic regexes help people build accurate mental models of LLM features.
@@ 2 Related Work
Interpretability aims to expose LLMs’ internal concepts in order to explain their behavior and evaluate their human-alignment (Lipton, 2018; Bricken et al., 2023; Olah et al., 2020; 2017). LLMs are known to encode concepts along linear directions in latent space, often referred to as features (or latents) (Park et al., 2024; Elhage et al., 2022). Methods like sparse autoencoders (SAEs) (Huben et al., 2024; Gao et al., 2025; Rajamanoharan et al., 2024; Bussmann et al., 2025) and transcoders (Ameisen et al., 2025; Dunefsky et al., 2024) extract LLM features, and have uncovered human-interpretable features, including those that correspond to domain-specific and safety relevant concepts (Lindsey et al., 2025; Gujral et al., 2025; Bricken et al., 2023).
While features correspond to concepts, which concept a feature encodes is not obvious a priori. Early interpretability work approached this problem by manually analyzing the inputs that activate each feature (Olah et al., 2020; 2017; 2018; Carter et al., 2019). To scale this process, automated interpretability aims to automatically describe each feature’s concept (Bau et al., 2017; Hernandez et al., 2022; Bills et al., 2023; Shaham et al., 2024). Given a feature’s activating data, automated interpretability prompts a language model describe the feature (Bills et al., 2023; Paulo et al., 2024; Gur-Arieh et al., 2025; Lin, 2023). Aligned with this goal, our method adopts a similar pipeline for generating feature descriptions, but generates descriptions using a structured language.
Our inspiration for describing LLM features with structured language follows a growing body of work that has found structure in the concepts features represent. Activation pattern analysis has revealed that feature complexity increases with depth (Jin et al., 2025) and that models represent concepts across levels of abstraction (Boggust et al., 2025; Chanin et al., 2024; Bussmann et al., 2025; Zaigrajew et al., 2025). In response, researchers have proposed lightweight taxonomies that classify features by their function,  (Ameisen et al., 2025; Lindsey et al., 2025; Gur-Arieh et al., 2025) and explicitly called for feature description formalisms (Huang et al., 2023). These findings motivate our use of a structured language, enabling semantic regexes to describe a feature’s activation pattern and encode its level of abstraction.
@@ 3 Semantic Regexes
Semantic regexes use a structured language to describe LLM features. Built around a system of human-interpretable primitives (Section 3.1.1) and modifiers (Section 3.1.2), semantic regexes are designed to describe the low-level syntactic patterns and higher-level semantic concepts that LLM features represent. Unlike natural language, which is flexible but ambiguous, the semantic regex language restricts expressivity to ensure the resulting feature descriptions explicitly convey their meaning. On the other hand, while semantic regexes are inspired by the formalism of regular expressions, they extend beyond one-to-one character patterns to capture more abstract concepts.
@@ 3.1.1 Primitives
Primitives are the atomic components of a sematic regex. They specify the type of textual pattern a semantic regex will match, ranging from exact characters to categorical relationships.
@@
Symbols ( [:symbol X:] ) match exact strings X . For instance, [:symbol color:] matches the string color, such as in hga hgsplash hgof hgcolor. These are the most specific and simplest primitives, and they commonly describe features that activate on specific tokens or phrases.
@@
Lexemes ( [:lexeme X:] ) match syntactic variants of X . Drawing on linguistics, a lexeme is the abstract form of a word that encompasses all of its surface variants, like changes in tense or plurality. For example, [:lexeme color:] matches color, colors, coloring, etc., such as hgcolor hgin hga hgcoloring hgbook. Lexemes typically describe features that capture a word’s meaning.
@@
Fields ( [:field X:] ) match semantic variants of X . Drawing on linguistics, fields refer to words or phrases in a conceptual domain. For instance, [:field color:] matches red, orange, blue, etc., as in hgblue hgjeans. Fields often apply to features that activate on a conceptual category.
@@ 4 Methods
To study how the semantic regex language alters feature interpretability, we embed it within a standard automated interpretability pipeline (Paulo et al., 2024; Templeton et al., 2024; Puri et al., 2025; Bills et al., 2023). The pipeline consists of three components: a subject model whose features we describe, an explainer model that generates feature descriptions, and an evaluator model that scores them. Given a subject model feature and its activating data, we prompt an explainer model to produce a description in natural or semantic regex language and use the evaluator model to score how well the description matches the feature’s behavior. This process allows us to directly compare semantic regexes to natural language descriptions across varied axes of interpretability.
@@ 4.1 Collecting Model Features and Activations
We study semantic regexes on two families of subject models, GPT-2 (Radford et al., 2019) and Gemma-2 (Mesnard et al., 2024). While our approach could describe model neurons (or any component matched to activating data), we focus on describing latent features identified by sparse coding methods (e.g., SAEs), since they often represent monosemantic concepts that are easier to describe and interpret (Bricken et al., 2023).
@@ 4.2 Generating Feature Descriptions
Given features and their activating data, we prompt an explainer model (gpt-4o-mini) to generate feature descriptions. We compare our semantic-regex method against natural language feature description methods: token-act-pair (Bills et al., 2023) and max-acts (Paulo et al., 2024). These methods are commonly used and vary in how they generate descriptions, allowing us to get a comprehensive understanding of how semantic regexes compare to natural language descriptions. See Section C.1 for implementation details.
@@
follows Neuronpedia’s (Lin, 2023) implementation of Paulo et al. (2024) (eleuther_acts_top2). Instead of showing each token’s activation value, like in token-act-pair, it maintains a more natural text format by delimiting activating tokens within the text, like “⟨⟨these tokens⟩⟩ activate”. It prompts the explainer model with the feature description task and three few-shot examples. Then, given the feature’s top 20 activating examples, the explainer model is asked to “describe the text latents that are common in the examples”.
@@
aims to accurately describe LLM features and only differs from natural language methods in the language available to the explainer model. As a result, we simply inject semantic-regex specific instructions into existing prompting strategies. To take advantage of its efficient example formatting, we adapt max-acts’ prompt by updating the instructions to follow the semantic regex language, adding a concise definition of the semantic regex language, and changing the few-shot examples to demonstrate semantic regex primitives and modifiers. We show the explainer model a feature’s top 10 activating examples and ask it to “output a short explanation followed by a semantic regex”, which we find improves the model’s ability to follow the syntax.
@@ 4.3 Evaluating Feature Descriptions
We evaluate feature descriptions using common automated interpretability metrics (Paulo et al., 2024; Puri et al., 2025). Each metric uses an evaluator model (gpt-4o-mini) to evaluate the descriptions’ fidelity to the feature’s behavior. Generation metrics test the description’s ability to generate activating examples (akin to precision), discrimination metrics test the description’s ability to match known activating examples (akin to recall), and faithfulness metrics test the description’s ability to match steered generation (a measure of causality). Implementation details in Section C.2).
@@
test a feature description’s precision by evaluating its ability to generate activating examples. We use clarity (Puri et al., 2025), which compares generated and random examples’ activations using the Gini index (a rescaling of the ROC AUC). Under this metric, overly broad descriptions score low by generating data outside the feature’s activation space, while ideal or overly narrow descriptions score high by only generating activating data.
@@
test the description’s recall by evaluating its ability to match known activating examples. We compute detection and fuzzing from Paulo et al. (2024) and responsiveness and purity from Puri et al. (2025). These metrics ask the evaluator model whether the feature description matches activating and random examples. Given the match results, detection measures balanced accuracy, responsiveness the Gini index, and purity average precision. Instead of matching the entire example, fuzzing asks if the description matches the example’s activating tokens and computes the balanced accuracy of these more specific match results. Under these metrics, overly narrow descriptions score low by missing activating examples, while ideal or overly broad descriptions score high by covering the entire activation space.
@@
test the description’s faithfulness to causal interventions on the feature. We use faithfulness (Puri et al., 2025), which asks the evaluation model whether the feature description matches continuations of random text when the feature is steered versus ablated.
@@ 5.1 Semantic Regexes are as Accurate as Natural Language Descriptions
We find that semantic-regex descriptions perform on par with natural language (Figure 2). Specifically, semantic regexes are non-inferior (p<0.05p<0.05222One-sided paired $t$-test with non-inferiority margin $\Delta=5\%$ and Bonferroni correction for superiority. ) to natural language on clarity across all models, detection for GPT-2-RES-25k, and responsiveness for Gemma-2-2B-RES-65k. Moreover, semantic-regex outperforms token-act-pair (p<0.05p<0.052) on clarity across all models, and detection, fuzzing, and responsiveness on GPT-2-RES-25k and Gemma-2-2B-RES-65k. This is non-obvious, as the semantic regex language is significantly constrained compared to the hundreds of thousands of words available in natural language. Moreover, while the explanation and evaluation models are well-versed in natural language, they learned the semantic regex language via only a brief description and few-shot examples. These results suggest that imposing structure on feature descriptions does not reduce their accuracy, while offering advantages that we explore in the following sections.
@@ 5.2 Semantic Regexes Improve Conciseness and Consistency
Concise feature descriptions adhere to explanation norms (Miller, 2019), making them easier to scan and interpret. While natural language descriptions often require verbose phrases to capture a feature’s activation pattern (Huang et al., 2023), semantic regexes can encode the same information in more compact form using its signal-rich components. For example, in Figure 3 top, the verbose natural language description “The presence of the sequence 54 indicating a year, time, or numeric reference frequently associated with events” can be expressed as [:symbol 54:] , and “Frequent use of verbs related to journalism that denote the act of covering or reporting on events or issues, often in a critical context” reduces to @{:context journalism:}([:lexeme cover:]) . Quantitatively, semantic regexes are consistently shorter. Across 100 randomly sampled features per layer in GPT-2-RES-25k, Gemma-2-2B-RES-16k, and Gemma-2-2B-RES-65k, the median description length is 41 characters (IQR: 19–59) compared to 139 (IQR: 119–166) for max-acts and 55 (IQR: 46–66) for token-act-pair. This compactness directly benefits interpretability, allowing human and model evaluators to parse shorter feature descriptions without sacrificing accuracy.

### v2

@@ Abstract
Automated interpretability aims to translate large language model (LLM) features into human understandable descriptions. However, natural language feature descriptions can be vague, inconsistent, and require manual relabeling. In response, we introduce semantic regexes, structured language descriptions of LLM features. By combining primitives that capture linguistic and semantic patterns with modifiers for contextualization, composition, and quantification, semantic regexes produce precise and expressive feature descriptions. Across quantitative benchmarks and qualitative analyses, semantic regexes match the accuracy of natural language while yielding more concise and consistent feature descriptions. Their inherent structure affords new types of analyses, including quantifying feature complexity across layers, scaling automated interpretability from insights into individual features to model-wide patterns. Finally, in user studies, we find that semantic regexes help people build accurate mental models of LLM features.
@@ 1 Introduction
Large language models (LLMs) represent their learned concepts, like the “Golden Gate Bridge”, as linear directions in latent space called features (Bricken et al., 2023). Understanding a model’s features helps us anticipate its behavior, assess its alignment, and intervene to ensure safe outcomes (Templeton et al., 2024). Automated interpretability assigns human-readable descriptions to each feature by analyzing patterns in its response to input data (Bills et al., 2023; Lin, 2023; Paulo et al., 2024). With feature descriptions, researchers have identified how models encode domain-relevant concepts, like protein structures (Gujral et al., 2025), and reconstructed circuits of features that correspond to complex behaviors, like medical diagnoses (Lindsey et al., 2025).
@@
In contrast to natural language, structured languages (e.g., regular languages, programming languages) offer well-defined syntax and semantics (Chomsky, 1956). By combining a set of primitives with compositional rules, structured languages capture precise patterns while maintaining expressivity (Lawson, 2005). Their grammatical structure provides additional affordances, including consistent ways of expressing the same pattern (Knuth, 1965), concise representations of complex patterns (Lawson, 2005), and mechanisms for representing abstraction (Backus, 1959; Aho & Ullman, 1972). As a result, they can construct complex yet specific expressions, ranging from regular expressions (Lawson, 2005) to software (Python Software Foundation, 2025).
@@
Across automated interpretability evaluations, we find that semantic regexes are as accurate as natural language, showing that constraining feature descriptions to the semantic regex language does not reduce expressivity. Beyond accuracy, their structured form offers additional affordances for interpretability, including producing more concise feature descriptions and enforcing description consistency across functionally similar features. Moreover, since semantic regex primitives and modifiers exist across levels of abstraction, they serve as a proxy for feature complexity, allowing us to perform model-wide analysis of feature behavior. Finally, in a user study, we find that semantic regexes help people build accurate mental models of LLM features.
@@ 2 Related Work
Mechanistic interpretability exposes LLMs’ internal concepts to explain their behavior and evaluate their alignment (Olah et al., 2017; 2020; Bricken et al., 2023). LLMs encode concepts along linear directions in latent space, often referred to as features (Elhage et al., 2022; Park et al., 2024). Methods like sparse autoencoders (SAEs) (Huben et al., 2024; Rajamanoharan et al., 2024; Bussmann et al., 2025; Gao et al., 2025) and transcoders (Dunefsky et al., 2024; Ameisen et al., 2025) uncover human-interpretable features, including those that correspond to domain-specific and safety-relevant concepts (Bricken et al., 2023; Gujral et al., 2025; Lindsey et al., 2025).
While features correspond to concepts, which concept a feature encodes is not obvious a priori. Early interpretability approached this problem by manually analyzing the data that activate each feature (Olah et al., 2017; 2018; 2020; Carter et al., 2019). To scale this process, automated interpretability prompts a language model to describe each feature based on its activating data (Bau et al., 2017; Hernandez et al., 2022; Bills et al., 2023; Lin, 2023; Shaham et al., 2024; Paulo et al., 2024; Gur-Arieh et al., 2025). Aligned with this goal, our method adopts a similar pipeline for generating feature descriptions, but generates descriptions using a structured language.
Our approach is inspired by recent research showing that LLM features represent structured concepts. Activation analyses show that feature complexity increases with depth (Jin et al., 2025) and models represent concepts across levels of abstraction (Chanin et al., 2024; Boggust et al., 2025; Bussmann et al., 2025; Zaigrajew et al., 2025). In response, researchers have proposed lightweight taxonomies that classify features by function (Ameisen et al., 2025; Lindsey et al., 2025; Gur-Arieh et al., 2025), and called for feature description formalisms (Huang et al., 2023). These findings motivate our use of a structured language, enabling semantic regexes to describe a feature’s activation pattern and express its level of abstraction.
@@ 3 Semantic Regexes
Semantic regexes use a structured language to describe LLM features. Built around a system of human-interpretable primitives (Section 3.1.1) and modifiers (Section 3.1.2), semantic regexes capture the low-level syntactic patterns and higher-level semantic concepts that LLM features represent. Unlike natural language, which is flexible but ambiguous, the semantic regex language restricts expressivity to ensure the resulting feature descriptions explicitly convey their meaning. On the other hand, while inspired by regular expressions, the semantic regex language is not a regular language and extends beyond one-to-one character patterns to capture more abstract concepts.
@@ 3.1 The Semantic Regex Language
We developed the semantic regex language using a grounded-theory approach (Corbin & Strauss, 1998), deriving its components from empirical analysis of real LLM features. By manually surveying thousands of features across models, layers, and feature sources on Neuronpedia (Lin, 2023), we identified recurring patterns (e.g., context dependent activations). We introduced new primitives or modifiers when they captured a recurring pattern, increased the language’s descriptive coverage, and preserved intelligibility of the language. We continued this iterative process until reaching saturation, resulting in a language capable of describing all the features we examined.
@@ 3.1.1 Primitives
Primitives are the atomic components of a semantic regex. They specify the type of textual pattern a semantic regex will match, ranging from exact characters to categorical relationships.
@@
Symbols ([:symbol X:]) match exact strings X. For instance, [:symbol color:] matches the string “color”, such as in hga hgsplash hgof hgcolor. These are the most specific and simplest primitives, and they commonly describe features that activate on specific tokens or phrases.
@@
Lexemes ([:lexeme X:]) match syntactic variants of X. Drawing on linguistics, a lexeme is the abstract form of a word that encompasses all of its surface variants, like changes in tense or plurality. For example, [:lexeme color:] matches “color”, “colors”, “coloring”, etc., such as hgcolor hgin hga hgcoloring hgbook. Lexemes typically describe features that capture a word’s meaning.
@@
Fields ([:field X:]) match semantic variants of X. Drawing on linguistics, fields refer to words or phrases in a conceptual domain. For instance, [:field color:] matches “red”, “blue”, etc., as in hgblue hgjeans. Fields often apply to features that activate on a conceptual category.
@@ 4 Methods
To study how the semantic regex language alters feature interpretability, we embed it within a standard automated interpretability pipeline (Bills et al., 2023; Paulo et al., 2024; Templeton et al., 2024; Puri et al., 2025). The pipeline consists of three components: a subject model whose features we describe, an explainer model that generates feature descriptions, and an evaluator model that scores descriptions. Given a subject model feature and its activating data, we prompt an explainer model to produce a description in natural or semantic regex language and use the evaluator model to score how well the description matches the feature’s behavior. Following prior interpretability work (Bills et al., 2023; Gur-Arieh et al., 2025) and our ablation study in Appendix C, we use GPT-4o-mini as the explainer and evaluator. This pipeline allows us to directly compare semantic regexes to natural language descriptions across varied axes of interpretability. It also demonstrates that by decoupling the format of the feature description from the generation process, semantic regexes are compatible with existing and future automated interpretability pipelines.
@@ 4.1 Collecting Model Features and Activations
We study semantic regexes on two families of subject models, GPT-2 (Radford et al., 2019) and Gemma-2 (Mesnard et al., 2024). While our approach could describe model neurons (or any component matched to activating data), we focus on describing latent features identified by sparse coding methods (e.g., SAEs) since they often represent monosemantic concepts that are easier to describe and interpret (Bricken et al., 2023).
@@ 4.2 Generating Feature Descriptions
Given features and their activating data, we prompt an explainer model (GPT-4o-mini) to generate feature descriptions. We compare our semantic-regex method against natural language feature description methods: token-act-pair (Bills et al., 2023) and max-acts (Paulo et al., 2024). These methods are commonly used and vary in how they generate descriptions, allowing us to get a broad understanding of how semantic regexes compare to natural language descriptions. See Section F.1 for implementation details.
@@
max-acts follows Neuronpedia’s (Lin, 2023) implementation of Paulo et al. (2024) (eleuther_acts_top2). Instead of showing each token’s activation value, like in token-act-pair, it maintains a more natural text format by delimiting activating tokens within the text, like “⟨⟨these tokens⟩⟩ activate”. It prompts the explainer model with the feature description task and three few-shot examples. Then, given the feature’s top 20 activating examples, the explainer model is asked to “describe the text latents that are common in the examples”.
@@
Since our goal is to accurately describe LLM features, semantic-regex only differs from natural language methods in the language available to the explainer model. We simply inject semantic regex specific instructions into existing prompting strategies. To take advantage of its efficient example formatting, we adapt max-acts’ prompt by updating the instructions to follow the semantic regex language, adding a concise definition of the semantic regex language, and changing the few-shot examples to demonstrate semantic regex primitives and modifiers. We show the explainer model a feature’s top 10 activating examples and ask it to “output a short explanation followed by a semantic regex”, which we find improves the model’s ability to follow the syntax.
@@ 4.3 Evaluating Feature Descriptions
We evaluate feature descriptions using common automated interpretability metrics (Paulo et al., 2024; Puri et al., 2025). Each metric uses an evaluator model (GPT-4o-mini) to evaluate the descriptions’ fidelity to the feature’s behavior. Generation metrics test the description’s ability to generate activating examples (akin to precision), discrimination metrics test the description’s ability to match known activating examples (akin to recall), and faithfulness metrics test the description’s ability to match steered generation (a measure of causality). Implementation details in Section F.2.
@@
Generation metrics evaluate a feature description’s ability to generate activating examples. We use clarity (Puri et al., 2025), which compares generated and random examples’ activations using the Gini index (a rescaling of the ROC AUC). Under this metric, overly broad descriptions score low by generating data outside the feature’s activation space, while ideal or slightly narrow descriptions score high by only generating activating data.
@@
Discrimination metrics test if the features description matches known activating examples. We compute detection and fuzzing from Paulo et al. (2024) and responsiveness and purity from Puri et al. (2025). These metrics ask the evaluator model whether the feature description matches activating and random examples. Given the match results, detection measures balanced accuracy, responsiveness the Gini index, and purity average precision. Instead of matching the entire example, fuzzing asks if the description matches the example’s activating tokens and computes the balanced accuracy of these more specific match results. Under these metrics, overly narrow descriptions score low by missing activating examples, while ideal or slightly broad descriptions score high by covering the entire activation space.
@@
Faithfulness tests how well the description reflects causal interventions on the feature. We use faithfulness (Puri et al., 2025), which asks the evaluator model whether the feature description matches continuations of random text when the feature is steered versus ablated.
@@ 5.1 Semantic Regexes are as Accurate as Natural Language Descriptions
Semantic regex descriptions perform on par with natural language (Figure 2). Specifically, semantic regexes are non-inferior (p<0.05p<0.05) [^1] One-sided paired $t$-test with non-inferiority margin $\Delta=5\%$ and Bonferroni correction for superiority. to natural language on clarity across all models, detection for GPT-2-RES-25k, and responsiveness for Gemma-2-2B-RES-65k. Moreover, semantic-regex outperforms token-act-pair (p<0.05p<0.05)1 on clarity across all models, and detection, fuzzing, and responsiveness on GPT-2-RES-25k and Gemma-2-2B-RES-65k. This is non-obvious, as the semantic regex language is significantly constrained compared to the grammatical flexibility of natural language. Moreover, while the explainer and evaluator models are well-versed in natural language, they learned the semantic regex language from only a brief description and few-shot examples. These results suggest that imposing structure on feature descriptions does not reduce their accuracy, while offering advantages that we explore in the following sections.
@@ 5.2 Semantic Regexes Improve Conciseness and Consistency
Concise feature descriptions adhere to explanation norms (Tim Miller, 2019), making them easier to scan and interpret. While natural language descriptions often require verbose phrases to capture a feature’s activation pattern (Huang et al., 2023), semantic regexes can encode the same information in a more compact form using its signal-rich components. For example, in Figure 3 top, the verbose natural language description “The presence of the sequence 54 indicating a year, time, or numeric reference frequently associated with events” can be expressed as [:symbol 54:]. Quantitatively, semantic regexes are consistently shorter. Across 100 randomly sampled features per layer in GPT-2-RES-25k, Gemma-2-2B-RES-16k, and Gemma-2-2B-RES-65k, the median description length is 41 characters (IQR: 19–59) compared to 139 (IQR: 119–166) for max-acts and 55 (IQR: 46–66) for token-act-pair. This conciseness directly benefits interpretability, allowing human evaluators in Section 5.4 to parse shorter descriptions without distracting details.